OpenStudy (anonymous):

m+6=3(1-m/2) please help i've been stuck on this for hours

jimthompson5910 (jim_thompson5910):

m+6=3(1-m/2) m+6=3(1)+3(-m/2) m+6=3-3m/2 Making sense so far?

OpenStudy (anonymous):

a little but not really :/

jimthompson5910 (jim_thompson5910):

Ok I'll post the full solution and you tell me where you get stuck (point out the line or step number) m+6=3(1-m/2) m+6=3(1)+3(-m/2) m+6=3-3m/2 2m+12=6-3m 2m=6-3m-12 2m+3m=6-12 5m=-6 m = -6/5

jimthompson5910 (jim_thompson5910):

I'm assuming of course that the equation is \[\Large m+6=3\left(1-\frac{m}{2}\right)\]

OpenStudy (anonymous):

yes thats it

jimthompson5910 (jim_thompson5910):

ok, then those steps above are correct

OpenStudy (anonymous):

i don't get why on the second step there are two 3's

jimthompson5910 (jim_thompson5910):

I distributed that 3 through

OpenStudy (anonymous):

over the 1-m/2?

jimthompson5910 (jim_thompson5910):

when you say something like 3(x+y) you distribute it through by multiplying 3 by x and 3 by y. Then you add like this 3x + 3y

OpenStudy (anonymous):

but there's only one m :/

jimthompson5910 (jim_thompson5910):

on the right side?

OpenStudy (anonymous):

oh wait there are two but it's over a fraction

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

i still am not understanding whatsoever i'm sorry :/

jimthompson5910 (jim_thompson5910):

that's ok, let me try it another way

OpenStudy (anonymous):

okay, i'm sorry

jimthompson5910 (jim_thompson5910):

no worries, don't be sorry

jimthompson5910 (jim_thompson5910):

does it look better if I write it like this \[\Large m+6=3\left(1-\frac{m}{2}\right)\]

jimthompson5910 (jim_thompson5910):

let me know

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok distribute to get \[\Large m+6=3\left(1\right)+3\left(-\frac{m}{2}\right)\] with me so far?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

that simplifies to \[\Large m+6=3-\frac{3m}{2}\]

OpenStudy (anonymous):

why does the 3 go automatically to the m? and not to the 2?

jimthompson5910 (jim_thompson5910):

\[\Large 3\left(-\frac{m}{2}\right)\] \[\Large \left(\frac{3}{1}\right)\left(-\frac{m}{2}\right)\] \[\Large \left(\frac{-3\times m}{1\times 2}\right)\] \[\Large -\frac{3m}{2}\] ------------------------------------------------------- So \[\Large 3\left(-\frac{m}{2}\right)\] simplifies to \[\Large -\frac{3m}{2}\]

OpenStudy (anonymous):

okay, now i understand

jimthompson5910 (jim_thompson5910):

so again, we have \[\Large m+6=3-\frac{3m}{2}\]

OpenStudy (anonymous):

yes :)

jimthompson5910 (jim_thompson5910):

With equations, you can clear out the fractions if you don't like them (as most legitimately do) I prefer to clear out fractions as well

jimthompson5910 (jim_thompson5910):

we do this by multiplying EVERY term by the LCD (which in this case is 2)

jimthompson5910 (jim_thompson5910):

\[\Large 2m+2*6=2*3+2\left(-\frac{3m}{2}\right)\] \[\Large 2m+12=6-\frac{6m}{2}\] \[\Large 2m+12=6-3m\]

jimthompson5910 (jim_thompson5910):

Notice how the fraction goes away after we do this

OpenStudy (anonymous):

i see

jimthompson5910 (jim_thompson5910):

Are you able to solve from here?

OpenStudy (anonymous):

no because i don't understand what to do getting 1 m over to the other m side

jimthompson5910 (jim_thompson5910):

You want to isolate m

jimthompson5910 (jim_thompson5910):

So you need to move all the 'm' terms to one side and get everything else to the other side

OpenStudy (anonymous):

but there are two m's

jimthompson5910 (jim_thompson5910):

so we need to get them together so we can combine them

jimthompson5910 (jim_thompson5910):

we do this by adding 3m to both sides

jimthompson5910 (jim_thompson5910):

\[\Large 2m+12=6-3m\] \[\Large 2m+12+3m=6-3m+3m\] \[\Large 5m+12=6\]

OpenStudy (anonymous):

okay! i understand now. elimination?

jimthompson5910 (jim_thompson5910):

now move the 12 over by subtracting 12 from both sides

jimthompson5910 (jim_thompson5910):

\[\Large 5m+12=6\] \[\Large 5m+12-12=6-12\] \[\Large 5m=-6\]

OpenStudy (anonymous):

and then divide both by -6?

jimthompson5910 (jim_thompson5910):

by 5

OpenStudy (anonymous):

or 5

jimthompson5910 (jim_thompson5910):

you want to move the 5 over, not the -6

OpenStudy (anonymous):

okay! so that would be -6/5????

jimthompson5910 (jim_thompson5910):

yes it would be

OpenStudy (anonymous):

thank you so much! :D i've been so stressed about this

jimthompson5910 (jim_thompson5910):

I'm glad to be of help